# Multi-Level Monte Carlo Training of Neural Operators

### v1

@@ 3.2 Message passing neural PDE solver
In this experiment, we consider the message-passing neural PDE solver architecture introduced by Brandstetter et al. (2022). This is a type of graph neural network (Kipf and Welling, 2016; Battaglia et al., 2018) written in the encoder-processor-decoder framework following earlier works (Battaglia et al., 2018; Sanchez-Gonzalez et al., 2020). We consider the problem of learning the time-forward velocity map $u(\mathbf{x},t)\mapsto u(\mathbf{x},t+\Delta t)$ associated with the Navier-Stokes equations on a classical fluid flow past a cylinder problem, near the Hopf bifurcation at Reynolds number $\textrm{Re}=200$, which features von Kármán vortex street patterns (see Section B.4). We generate four meshes $M_{4},\ldots,M_{1}$ with decreasing mesh resolution using the Gmsh meshing software (Geuzaine and Remacle, 2009), and solve the corresponding time-dependent equations by means of finite elements and Crank-Nicolson time-stepping scheme using the Firedrake library (Rathgeber et al., 2016) on each mesh. This contrasts with previous experiments of Section 3.1 (and what is usually done in operator learning), where the numerical solutions are acquired on the finest resolution and down-sampled to coarser ones. Here, we aim to evaluate the performance of MLMC training in the challenging settings where the coarse data quality is low (since they are generated by a numerical solver on a coarser spatial discretization).
Table 1: Fluid flow past cylinder problem. Baselines average training time per epoch and test loss when training the graph neural network for varying resolutions $R_{i}$, along with the results using MLMC training. Best trade-off, defined as training time multiplied by test MSE, is in bold, while second best is underlined.
@@
MSE test loss [×$\times 10^{-2}$] 5.0 5.3 5.7 0.5 1.1 0.8
@@ 3.3 Geometry-informed neural operator transformer
GINOT  (Liu et al., 2025) is a neural operator architecture based on transformer networks Vaswani et al. (2017) that is designed to learn mappings between complex geometries (represented as CAD meshes or cloud of points) to corresponding solution fields (see Fig. 4). It consists of two main components: a geometry encoder and a solution decoder, which are combined by the cross-attention mechanism of the transformer network.
@@
In this numerical experiment, we apply the MLMC neural operator training algorithm to the GINOT architecture on the complex DeepJEB dataset (Hong et al., 2024) containing 3D large-scale jet engine brackets and corresponding von Mises stress fields under vertical loading. The original dataset consists of sample geometries with resolutions ranging from $R_{\text{test}}\in[18085,55009]$, referred to as the testing resolution (as used in Liu et al. (2025)). To enable MLMC training, we downsample the data to $R_{2}\in[4521,13752]$ and $R_{1}\in[282,859]$, which we use as the fine and coarse resolutions, respectively. Further details on the dataset are available in Section B.5. Sample geometries from the dataset at different resolutions are depicted in Fig. 4(a). We train the GINOT on the DeepJEB dataset at different resolutions $R_{1}$, $R_{2}$, and $R_{\text{test}}$ over $500$ epochs and report the average training time per epoch along with the test loss on the resolution $R_{\text{test}}$ in Table 2. We note that we slightly modified the GINOT architecture and were able to outperform the testing error of $35.6\%$ on the resolution $R_{\text{test}}$ reported by Liu et al. (2025).
@@
In the last row of Table 2, we show the performance (on the finest resolution $R_{\text{test}}$) of the MLMC training algorithm with the two levels of resolution $R_{1}$ and $R_{2}$, together with with the corresponding computational training time per epoch. In this complex task, the proposed MLMC approach achieves a testing loss of $26.3\%$ with a training time of $20.1$ seconds per epoch, which improves upon the Pareto frontier of the training time/accuracy trade-off, between the coarse ($R_{1}$) and fine ($R_{2}$) resolution training baselines. We note that the training time is nearly halved compared to the fine resolution training without loosing too much accuracy. In Fig. 4(b), we display one ground truth sample at resolution $R_{\text{test}}$ and its MLMC approximation obtained by considering two levels of resolutions $m=2$, with a nested sub-sampling as pairing strategy, $1723$ and $213$ samples at level one and two (optimal sample progression), and a batch multiplier of $\delta=2$. Finally, in Fig. 5, we display the corresponding approximation of the fine gradients as we include the fine resolution correction term in the telescopic sum of Eq. 3.
@@ 4 Conclusions
[…]
Current limitations, that can be addressed by future developments, are related to the need of sufficiently complex model architecture in order to benefit from the multi-resolution approach. The optimal sampling distribution depends on problem-specific variance decay rates that may need fine-tuning for each application. Finally, approximation errors from coarse representations might be dominating in problems with extremely sharp features or discontinuities, and potentially reduce the efficiency of the MLMC approach. Additional future developments could include dynamically adjusting the sampling distributions across resolutions during training. Finally, MLMC is a flexible approach, which leads itself to pretraining and finetuning applications, and possibly extensions to online learning scenarios, thus enabling efficient adaptation of pre-trained models to new data distributions with minimal computational overhead.
Acknowledgments and Disclosure of Funding

### v2

@@ 3.2 Message passing neural PDE solver
In this experiment, we consider the message-passing neural PDE solver architecture introduced by [5]. This is a type of graph neural network [22, 2] written in the encoder-processor-decoder framework following earlier works [2, 39]. We consider the problem of learning the time-forward velocity map $u(\mathbf{x},t)\mapsto u(\mathbf{x},t+\Delta t)$ associated with the Navier-Stokes equations on a classical fluid flow past a cylinder problem, near the Hopf bifurcation at Reynolds number $\textrm{Re}=200$, which features von Kármán vortex street patterns. This is a well-known benchmark problem for the evaluation of numerical algorithms for incompressible time-dependent Navier-Stokes equations [40]. The equations read as
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}-\nabla\cdot\frac{2}{\textnormal{Re}}\boldsymbol{\epsilon}(\mathbf{u})+\nabla p$ $=0,$ $\text{in }\Omega\times(0,T],$ (9)
$\nabla\cdot\mathbf{u}$ $=0,$ $\text{in }\Omega\times(0,T],$
$\mathbf{u}(\mathbf{x},0)$ $=\mathbf{u}_{0}(\mathbf{x}),$ $\mathbf{x}\in\Omega,$
where $\Omega\subset\mathbb{R}^{2}$, $T=50$, $p:\Omega\times[0,T)\rightarrow\mathbb{R}$ is the pressure field, the Reynolds number is Re = 200, and $\boldsymbol{\epsilon}(\mathbf{u})=(\nabla\mathbf{u}+\nabla\mathbf{u}^{\top})/2$. The boundary conditions are defined as follows: $\mathbf{u}=(1,0)^{\top}$ on the upper and lower side of $\Omega$, a homogeneous Dirichlet boundary condition on the disk obstacle, and a “do-nothing” condition on the right side of the domain. We consider fluid fields including laminar flow as well as vortex structures. The dataset is formed of $N=200$ samples obtained by considering different random inflow conditions on the left boundary, which are split as $150$ for training and the remaining $50$ for testing. The equations are discretized in space using standard Taylor-Hood finite elements for velocity and pressure, which is a stable element pair for the Stokes equation [41], and in time by a Crank–Nicolson time-stepping scheme with a timestep of $0.25$. The velocity and pressure solutions are sampled every $\Delta t=1$ timesteps, and we aim to learn the time-forward operator $\mathbf{u}(\cdot,t)\mapsto\mathbf{u}(\cdot,t+\Delta t)$, by mapping the first $30$ time steps to the next $20$. The finest mesh $M_{4}$ on which the simulations are performed contains $14631$ triangle elements with $7217$ nodes, and a target element size of $0.03$ near the obstacle and $0.12$ on the boundary of the domain. The computed solutions are projected onto a space of linear finite elements and the values at the mesh nodes are saved. We also generated coarser meshes using the Gmsh meshing software [12] with approximately twice fewer nodes from level $i+1$ to level $i$ (specifications are available in Table 1) to run the MLMC algorithm, and performed numerical simulations on these coarser meshes to acquire the corresponding data. This contrasts with previous experiments of Section 3.1 (and what is usually done in operator learning), where the numerical solutions are acquired on the finest resolution and down-sampled to coarser ones. Here, we aim to evaluate the performance of MLMC training in the challenging settings where the coarse data quality is low (since they are generated by a numerical solver on a coarser spatial discretization).
Mesh name #\# nodes #\# elements mesh size [obstacle] mesh size [boundary]
$M_{4}$ 7217 14631 0.05 0.18
$M_{3}$ 3808 7760 0.07 0.25
$M_{2}$ 1645 3387 0.10 0.4
$M_{1}$ 806 1682 0.14 0.6
Table 2: Fluid flow past cylinder problem. Baselines average training time per epoch and test loss when training the graph neural network for varying resolutions $M_{i}$, along with the results using MLMC training. Best trade-off, defined as training time multiplied by test MSE, is in bold, while second best is underlined.
@@
MSE test loss $\times 10^{-2}$] 5.0 5.3 5.7 0.5 1.1 0.8
3.3 Mesh-free convolutional neural network
The architecture considered is a resolution-invariant CNN specifically designed for mesh-based inputs, introduced in [44, 38]. The main idea is to define convolution operations without relying on a mesh of the domain $\Omega\subset\mathbb{R}^{2}$, adopting a mesh-free approach instead. More precisely, the mesh-free convolution reads as
The elliptic differential operator in Eq. 11 is defined as
$\mathcal{L}(\boldsymbol{\alpha})\;=\;\alpha_{xx}\,\frac{\partial^{2}v}{\partial x^{2}}\;+\;2\,\alpha_{xy}\,\frac{\partial^{2}v}{\partial x\,\partial y}\;+\;\alpha_{yy}\,\frac{\partial^{2}v}{\partial y^{2}},$
where $\alpha_{xx},\alpha_{yy}>0$ and $\alpha_{xx}\alpha_{yy}-\alpha_{xy}^{2}>0$ thus ensuring that $\mathcal{L}$ is elliptic and invertible. The vector $\boldsymbol{\xi}=[\tau,\boldsymbol{\alpha},\boldsymbol{\gamma}]\in\mathbb{R}^{7}$ collects all the trainable parameters. The input field $a$ is mapped to the embedding $\tilde{v}$ by means of the operator $\mathcal{R}:\mathcal{A}\rightarrow\mathbb{R}^{d_{v}}$ consisting of a convolution operation Eq. 10 with $1\times 1$ filters. In this way, the initial condition of Eq. 11 is obtained and the parabolic PDE is solved. Finally, the output of the convolution is obtained by taking the derivative of $\tilde{v}(\tau,\mathbf{x})$ in the direction $\gamma$. So, introducing a suitable uniform discretization of the domain $\Omega$, solving Eq. 11 by means of Fast Fourier Transform (FFT) and differentiating the solution in the direction $\boldsymbol{\gamma}$ actually consists of applying the mesh-free convolution [44].
Advection-diffusion equation
$u(\mathbf{x})$ $=0,$ $\mathbf{x}\in\Gamma_{D},$
$\nu\frac{\partial u}{\partial n}$ $=0,$ $\mathbf{x}\in\Gamma_{N}.$
This equation models the transport and diffusion of the pollutant released by industrial chimneys into a plane [9]. Here, $\mathbf{b}=(0.0025,0.0025)^{\top}$, $f(\mathbf{x})=\sum_{i=1}^{4}f_{i}\chi_{D_{i}}(\mathbf{x})$, where $f_{i}\geq 0$, and $D_{i}$ the spatial region representing the $i$-th chimney and $f_{i}$ its emission rate (normalized with respect to an appropriate emission volume). We impose a homogeneous Dirichlet condition over the inflow boundary $\Gamma_{D}=\{\mathbf{x}\in\partial\Omega:\mathbf{b}\cdot\mathbf{n}(\mathbf{x})<0\}$, where $\mathbf{n}(\mathbf{x})$ is the unit vector directed outward, and homogeneous Neumann condition on the outflow boundary $\Gamma_{N}=\partial\Omega\setminus\Gamma_{D}$. For the expression of the diffusion field $\nu$ we refer to [9]. Eq. 12 is discretized in space using linear finite elements by considering $\{R_{i}\times R_{i}\}_{i=1}^{3}=\{75\times 75,150\times 150,300\times 300\}$ space points. The dataset consists of $N=1024$ samples obtained by drawing instances of $x_{c}^{i}\sim\mathcal{U}(-3,-2)$ and $r^{i}\sim\mathcal{U}(0.1,0.3)$, parametrizing the regions $D_{i}$, for $i\in\{1,\ldots,4\}$.
We validate the proposed algorithm by employing mesh-free convolutional neural networks on the test cases Eqs. 7 and 12. In particular, to take into account advection and incorporate the knowledge we have about the system in Eq. 12, we add to the mesh-free convolutional kernel a term of the form [38]:
$\psi_{x}[-1,0,1]\quad\textnormal{and}\quad\psi_{y}[1,0,-1],$
that is we add a second convolutional layer by fixing their weights equal to the second order finite difference discretization of the advection term, and we multiply the output by a trainable parameter in each direction of the domain. Then, the vector of all trainable parameters becomes $\boldsymbol{\xi}=[\tau,\boldsymbol{\alpha},\boldsymbol{\gamma},\boldsymbol{\psi}]\in\mathbb{R}^{9}$ for each mesh-free convolutional layer.
Table 3: Darcy problem. (a) Baselines training time per epoch and test loss by varying the resolution $R$. (b) MLMC training time per epoch and test loss by varying $\delta$ and $m=3$.
Resolution $R=15$ $R=30$ $R=60$ $R=120$ $R=241$
Train epoch time $[s]$ 0.81 0.83 0.91 1.26 3.96
Test loss $\times 10^{-2}$] 15.2 2.5 2.0 2.0 1.8
Train epoch time $[s]$ 0.19 0.55 1.25 1.66 2.48 0.60
Test loss $\times 10^{-2}$] 7.8 4.7 3.3 3.2 3.1 2.5
Test on Darcy problem
To obtain the baselines, we first run standard SGD to approximate the solution operator on the resolutions $\{R_{m}\}_{m=1}^{5}=\{241,120,60,30,15\}$ and the results are the ones reported in Table 3(a). The training time per epoch and the testing loss values obtained by means of our proposed training paradigm, with $m=3$ levels and hierarchy as pairing strategy, are gathered in Table 3(b). As expected, by increasing the multiplier $\delta$ the test loss increases as the epoch train time decreases because a smaller amount of fine training data is used. We also report the test related to the use of the optimal number of samples per resolution $N_{i}^{*}$ (Proposition 2.2 with $k=1$ and $d=2$). Moreover, we select $\delta=\delta^{*}=2$. This configuration is the one providing the best performance, in terms of efficiency, with respect to the baseline with $R=30$ (the baseline resolution providing a comparable accuracy), and it is related to the fact that even if $N_{i}^{*}\approx 2^{-2i}N_{i-1}^{*}$ we are employing $\delta=2$ instead of $\delta=8$; indeed, in the latter case, during one epoch less batches of fine resolution data are seen with respect to the number of batches of coarse data. As the value for $\delta$ increases the gap between the number of batches of fine and coarse data increases making the gradients’ estimates more and more biased towards large-scales, or in the same way, the exposure to fine-scale corrections is limited. In Fig. 9, we report the comparison among the ground truth data and the approximation obtained by means of traditional SGD, both with $R=241$, and the MLMC approximated solution, with hierarchy pairing strategy, $m^{*}=3$ and $\delta^{*}=2$, for different testing instances of the diffusion coefficients.
Test on advection-diffusion
[…]
@@ 3.4 Geometry-informed neural operator transformer
GINOT  [33] is a neural operator architecture based on transformer networks [42] that is designed to learn mappings between complex geometries (represented as CAD meshes or cloud of points) to corresponding solution fields (see Fig. 10). It consists of two main components: a geometry encoder and a solution decoder, which are combined by the cross-attention mechanism of the transformer network.
@@
In this numerical experiment, we apply the MLMC neural operator training algorithm to the GINOT architecture on the Deep jet engine bracket (DeepJEB) dataset [19] containing 3D large-scale jet engine brackets and corresponding von Mises stress fields under vertical loading. The DeepJEB datatest is publicly accessible and comprises $N=2138$ samples of complex 3D jet engine brackets with arbitrary shapes and sizes. More precisely, the dataset was created using deep generative models to perform data augmentation on the simulated jet engine bracket (SimJEB) dataset [43] thus enhancing the diversity of shapes and the performance features by means of AI. SimJEB is one of the largest publicly available dataset for jet engine bracket, and includes 381 hand-designed CAD models, along with finite element simulation results of von Mises stress and displacement fields under different loading conditions. Following [33], we focus on the approximation of the von Mises stress data under vertical loading. The DeepJEB dataset volume mesh nodes ranges from $18085$ to $55009$, which we refer to as testing resolution $R_{\text{test}}$ in this paper. We use it to evaluate the super-resolution capabilities of the trained transformer model [33]. To apply the MLMC training algorithm, we generate a similar dataset at a coarse level of resolution $R_{1}\in[282,859]$ and a fine level of resolution $R_{2}\in[4521,13752]$ by downsampling the original dataset. The dataset downsampling was performed using the PointNet++ [37] algorithm exploiting the pool.fps command from the torch_geometric library [10] with a default sampling ratio of $0.5$.
@@
We consider the same hyperparameters as in [33] for the GINOT architecture, except for the number of attention heads in the encoder and decoder, which we set to $4$ instead of $1$, encoder/decoder cross-attention layers (set to $4$) and encoder self-attention layers (set to $4$). By using a larger architecture, we are able to improve the performance of the GINOT model on the test resolution and reach a normalized mean square error of $18.6\%$, which is much lower than $35.6\%$ reported in [33]. Sample geometries from the dataset at different resolutions are depicted in Fig. 10(a). We train the GINOT on the DeepJEB dataset at different resolutions $R_{1}$, $R_{2}$, and $R_{\text{test}}$ over $500$ epochs and report the average training time per epoch along with the test loss on the resolution $R_{\text{test}}$ in Table 4. We note that we slightly modified the GINOT architecture and were able to outperform the testing error of $35.6\%$ on the resolution $R_{\text{test}}$ reported by [33].
@@
[…]
@@ 4 Conclusions
Neural operators have shown remarkable promise in the approximation of operators associated with PDEs, but their practical adoption is hindered by the computational cost of high-resolution training. We introduced a mathematically rigorous MLMC approach to efficiently train neural operators by approximating fine-resolution gradients through a telescopic sum formulation across mesh hierarchies or function discretizations. Our implementation is characterized by a careful design of resolution-aware data structures and optimized memory management. Our numerical results, across a variety of state-of-the-art neural operator architectures and complex benchmarks, demonstrated that the proposed MLMC training algorithm accelerates training while maintaining the same accuracy.
Current limitations, that can be addressed by future developments, are related to the need for sufficiently complex model architecture in order to benefit from the multi-resolution approach. The optimal sampling distribution depends on problem-specific variance decay rates that may need fine-tuning for each application. Finally, approximation errors from coarse representations might be dominating in problems with extremely sharp features or discontinuities, and potentially reduce the efficiency of the MLMC approach. Additional future developments could include dynamically adjusting the sampling distributions across resolutions during training. Finally, MLMC is a flexible approach, which leads itself to pretraining and finetuning applications, and possibly extensions to online learning scenarios, thus enabling efficient adaptation of pretrained models to new data distributions with minimal computational overhead.
Acknowledgements